OpenStudy (anonymous):

What is the solution to the equation log base 2 of x – log base 2 of 4 = 5 ?

OpenStudy (mertsj):

logbase2 of x/4 = 5 So 2^5 = x/4 32 = x/4 x = 4(32)=128